(1 point) Consider the following series. Σε 25 Find the first five partial sums of the series. S2 = 2/8 0.4166666
Answer #1

2 = 0.25 s. = 2 = S2 2C742) 2 (2+2 ) S = + is $ = 2212 + 36 3 (372) 2 (2+2) = = 0.25 +0.13333 0.30333 tw S4 = meninta) n(nt.Therefore 2 = 2/81 = 10:25] = و = وی 0 . 3833 0 . 4666 = 505231 ا 005653 = و 2020 -3- 12 03:45